solve the systems of equations by graphing
-3x-2y=6
3x-y=12

graph and give a solution set?

Found 2 solutions by fractalier, FrankM:
Answer by fractalier(6550)

Add these together and get
-3x-2y=6
3x-y=12
--------------
-3y = 18
y = -6
so that
-3x -2(-6) = 6
-3x + 12 = 6
-3x = -6
x = 2
(2, -6)

Answer by FrankM(1040)